# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v1

@@ B.1. Proof of Theorem 3
(68) $G^{\star}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{G^{\epsilon}_{n}(y)\,\mid\,|x-y|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\}.$
Thus, in passing from $F$’s to $G$’s, nothing has changed except the tighter upper bound $G^{\epsilon}_{n}\leq c_{+}-\delta<1$ holds for each $n$ and the evolution equation has been replaced by an inequality in (67).
Suppose that $\varphi$ is a smooth function defined in some open subset of space-time that touches $G^{\star}$ from above at some point $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ in its domain. The goal is to show that $\partial_{t}\varphi(x_{0},t_{0})\leq\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))$. Notice that one technicality can be ruled out immediately: Since $G^{\star}\geq c_{-}$ pointwise, if $\varphi(x_{0},t_{0})=c_{-}$, then
$\partial_{t}\varphi(x_{0},t_{0})=\partial_{x}\varphi(x_{0},t_{0})=0,\quad\partial_{x}^{2}\varphi(x_{0},t_{0})\geq 0,$
which implies the desired inequality by (51). Therefore, from now on, assume that $c_{-}<\varphi(x_{0},t_{0})$.
According to Lemma 3, there is no loss of generality in assuming that either (a) $|\partial_{x}\varphi(x_{0},t_{0})|>0$ or (b) $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$. In case (a), since $G^{\star}$ is nondecreasing, it follows that $\partial_{x}\varphi(x_{0},t_{0})>0$, and Lemma 1 is applicable. In case (b), Lemma 2 applies. To avoid repetition, only case (b) will be considered in what follows.
Therefore, assume henceforth that (b) holds. According to Lemma 2 with $A=c_{+}-\frac{\delta}{2}$, there is no loss of generality in fixing an $r>0$ and assuming that $\varphi(x,t)=\varphi_{1}(x)+\varphi_{2}(t)$ for $(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]$, where $\varphi_{1}$ is a nondecreasing $C^{2}$ function with $\varphi_{1}(x)=\varphi_{1}(x_{0})$ for $x\leq x_{0}$ and $\varphi_{1}(x)=c_{+}-\frac{\delta}{2}$ for $x\geq x_{0}+r$; $\varphi_{2}$ is $C^{2}$ in $t_{0}-r,t_{0}+r]$ with $\varphi_{2}(t_{0})=0$; and
@@
Since $\varphi_{2}$ is continuous in $t_{0}-r,t_{0}+r]$ and $\varphi_{1}(-\infty)=\varphi(x_{0},t_{0})>c_{-}$, up to shrinking $r$ if necessary, there is no loss of generality in assuming that there are constants $c_{*},c^{*}\in[0,1]$ such that, for each $t\in[t_{0}-r,t_{0}+r]$,
(69) $c_{-}<c_{*}\leq\varphi(-\infty,t)\quad\text{and}\quad c_{+}-\delta<c_{*}\leq\varphi(+\infty,t)\leq c_{+}-\frac{\delta}{4}.$
Finally, for each $\epsilon>0$, let $c(\epsilon)$ be the constant
$c(\epsilon)=\sup\left\{G^{\epsilon}_{n}(x)-\varphi(x,\epsilon^{\alpha}n)\,\mid\,x\in\mathbb{R},\,\,n\in\mathbb{N}\cap\epsilon^{-\alpha}[t_{0}-r,t_{0}+r]\right\}.$
Notice that (69) and (66) imply that there is an $R>0$ such that, for any $\epsilon>0$,
$\sup\left\{G^{\epsilon}_{n}(x)-\varphi(x,\epsilon^{\alpha}n)\,\mid\,|x|\geq R,\,\,n\in\mathbb{N}\cap\epsilon^{-\alpha}[t_{0}-r,t_{0}+r]\right\}<0.$
Therefore, by the equation (68) relating $G^{\epsilon}_{n}$ to $G^{\star}$, it follows that there is a sequence $\{\epsilon_{k},x_{k},n_{k}\}$ such that $(\epsilon_{k},c(\epsilon_{k}))\to(0,0)$ as $k\to\infty$; $|x_{k}|\leq R$ and $|\epsilon_{k}^{\alpha}n_{k}-t|\leq r$ for each $k$; and
(70) $G^{\epsilon_{k}}_{n_{k}}(x_{k})-\varphi(x_{k},\epsilon_{k}^{\alpha}n_{k})-c(\epsilon_{k})=0.$
Further, since $\{(x,t)\,\mid\,G^{\star}(x,t)=\varphi(x,t)\}$ is contained in $(-\infty,x_{0}]\times\{t_{0}\}$, there is no loss of generality in assuming that there is a $y_{0}\leq x_{0}$ such that $(x_{k},\epsilon^{\alpha}_{k}n_{k})\to(y_{0},t_{0})$ and $G^{\star}(y_{0},t_{0})=\varphi(y_{0},t_{0})$. Note, in particular, that the inequality $y_{0}\leq x_{0}$ implies
(71) $\lim_{k\to\infty}(\partial_{x}\varphi_{1}(x_{k}),\partial_{x}^{2}\varphi_{1}(x_{k}))=(0,0).$
Finally, for notational ease, let $\Phi_{k}:\mathbb{R}\times[t_{0}-r,t_{0}+r]\to\mathbb{R}$ be the function $\Phi_{k}(x,t)=\varphi(x,t)+c(\epsilon_{k})$. Since $c(\epsilon_{k})\to 0$, the bounds (69) implies that $0\leq\Phi_{k}\leq 1$ pointwise for all $k$ large enough. That is, $\Phi_{k}\in CDF(\overline{\mathbb{R}})$ for large enough $k$. By the choice of $c(\epsilon_{k})$,
$G^{\epsilon_{k}}_{n}\leq\Phi_{k}(\cdot,\epsilon^{\alpha}_{k}n)\quad\text{pointwise for each}\,\,n\in\mathbb{N}\cap\epsilon_{k}^{-\alpha}[t_{0}-r,t_{0}+r].$
Therefore, by the monotonicity of $T_{\epsilon}$,
$T_{\epsilon_{k}}G^{\epsilon_{k}}_{n_{k}-1}\leq T_{\epsilon_{k}}\Phi_{\epsilon_{k}}(\cdot,\epsilon^{\alpha}_{k}n_{k}-\epsilon^{\alpha}_{k}).$
@@
$\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k})=G^{\epsilon_{k}}_{n_{k}}(x_{k})\leq T_{\epsilon_{k}}G^{\epsilon_{k}}_{n_{k}-1}(x_{k})\leq T_{\epsilon_{k}}\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon_{k}^{\alpha}).$
@@
$\epsilon_{k}^{-\alpha}\Big(\Phi_{\epsilon_{k}}(x_{k},\epsilon_{k}^{\alpha}n_{k})-\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon^{\alpha}_{k})\Big)-\epsilon_{k}^{-\alpha}(\mathscr{L}_{\epsilon_{k}}\Phi_{k})(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon_{k}^{\alpha})\leq 0.$
By assumption (ii),
$\mathscr{L}_{\epsilon_{k}}\Phi_{k}(x_{k},\epsilon^{\alpha}_{k}n_{k}-\epsilon_{k}^{\alpha})=\mathscr{L}_{\epsilon_{k}}\varphi_{1}(x_{k}),$
hence, by assumption (iv), (71), and the form of $\varphi$, sending $k\to\infty$ yields
$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(0,0)=\partial_{t}\varphi_{2}(t_{0})-\lim_{k\to\infty}\mathcal{F}(\partial_{x}\varphi_{1}(x_{k}),\partial_{x}^{2}\varphi_{1}(x_{k}))\leq 0.$
Since $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$ by assumption, this is the desired inequality.
Step 5: $F_{\star}$ is a Supersolution. The proof that $F_{\star}$ is a supersolution of (57) can be reduced to the subsolution property of $F^{\star}$ after a suitable transformation. As in Remark 2 above, let $G_{-}$ denote the left-continuous version of a nondecreasing function $G$, and define an involution $G\mapsto G_{\text{rev}}$ on $CDF(\overline{\mathbb{R}})$ via the formula
@@
Let $\{\tilde{T}_{\epsilon}\}$ be the operators on $CDF(\overline{\mathbb{R}})$ obtained from $\{T_{\epsilon}\}$ by
$\tilde{T}_{\epsilon}F=(T_{\epsilon}(F_{\text{rev}}))_{\text{rev}}.$
It is straightforward to check that the family $\{\tilde{T}_{\epsilon}\}$ satisfies assumptions (i)-(iv), the only difference being $\mathcal{F}$ has to be replaced by the function $\tilde{\mathcal{F}}(p,X)=-\mathcal{F}(p,-X)$.
Define $\{\tilde{F}^{\epsilon}_{n}\}$ by the rule $\tilde{F}^{\epsilon}_{n}=(F^{\epsilon}_{n})_{\text{rev}}$. Notice that the recursion $F^{\epsilon}_{n+1}=T_{\epsilon}F^{\epsilon}_{n}$ implies that $\tilde{F}^{\epsilon}_{n+1}=\tilde{T}_{\epsilon}\tilde{F}^{\epsilon}_{n}$. Therefore, if $\tilde{F}^{\star}$ is defined by
$\tilde{F}^{\star}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{\tilde{F}^{\epsilon}_{n}(y)\,\mid\,|y-x|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\},$
then $\tilde{F}^{\star}$ is a subsolution of the initial value problem (57) with $\mathcal{F}$ replaced by $\tilde{\mathcal{F}}$ and $F_{\text{in}}$ by $(F_{\text{in}})_{\text{rev}}$. Further, by definition, $\tilde{F}^{\star}(x,t)=1-F_{\star}(-x,t)$. From these last two observations, one readily concludes that $F_{\star}$ is a viscosity supersolution of (57).∎
@@ B.2. Proof of Corollary 3
In the proof that follows, it it convenient to work with the half-relaxed limits $\liminf_{*}F_{\epsilon}$ and $\limsup_{*}F_{\epsilon}$ of the rescaled CDF $F_{\epsilon}$ defined by
$(\liminf\nolimits_{*}F_{\epsilon})(x,t)$ $=\lim_{\delta\downarrow 0}\inf\left\{F_{\epsilon}(y,s)\,\mid\,|y-x|+|s-t|+\epsilon\leq\delta\right\},$
$(\limsup\nolimits_{*}F_{\epsilon})(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{\epsilon}(y,s)\,\mid\,|y-x|+|s-t|+\epsilon\leq\delta\right\}.$
@@
Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. Suppose that $F_{\epsilon}(x,0)\to F_{\text{in}}(x)$ at each point of continuity $x$ of $F_{\text{in}}$. Let $\overline{F}$ and $\underline{F}$ be the maximal subsolution and the minimal supersolution of the problem (57). In view of Proposition 9 and well-known properties of half-relaxed limits (see [13, Remark 6.4] or [7, Lemma 6.2]), to see that $F_{\epsilon}$ converges locally uniformly to the discontinuous viscosity solution, it suffices to establish that
$\overline{F}\leq\liminf\nolimits_{*}F_{\epsilon}\leq\limsup\nolimits^{*}F_{\epsilon}\leq\underline{F}\quad\text{in}\,\,\mathbb{R}\times(0,\infty).$
@@
For each $N\in\mathbb{N}$, define $F_{\epsilon,N}$ like $F_{\epsilon}$, albeit instead with the initial datum
$F_{\epsilon,N}(x,0)=2^{N}\int_{0}^{\infty}F_{\epsilon}(x-y,0)\rho(2^{N}y)\,dy,$
where $\rho:(0,\infty)\to(0,\infty)$ is a smooth function supported in $[1,2]$ with $\int_{0}^{\infty}\rho(y)\,dy=1$. From the fact that $\rho$ is supported in $[1,2]$ and $F_{\epsilon}$ is nondecreasing in $x$, it follows that $F_{\epsilon,N}(x,0)\leq F_{\epsilon,N+1}(x,0)\leq F_{\epsilon}(x,0)$ for every $x\in\mathbb{R}$. Due to the monotonicity assumption (i), this implies that $F_{\epsilon,N}\leq F_{\epsilon,N+1}\leq F_{\epsilon}$ in $\mathbb{R}\times[0,\infty)$ for each $\epsilon>0$ and $N\in\mathbb{N}$. By Theorem 3, $F_{\epsilon,N}\to F_{N}$ as $\epsilon\to 0$, where $F_{N}$ is the solution of (57) with initial datum $F_{\text{in}}^{(N)}$ given by $F_{\text{in}}^{(N)}(x)=2^{N}\int_{0}^{\infty}F_{\text{in}}(x-y)\rho(2^{N}y)\,dy$. Note that $F_{N}\leq\liminf\nolimits_{*}F_{\epsilon}$ pointwise in $\mathbb{R}\times[0,\infty)$ for any fixed $N$.
Standard arguments (see [7, Theorem 6.2]) show that the function $F_{\infty}$ defined by
$F_{\infty}(x,t)=\lim_{\delta\downarrow 0}\inf\left\{F_{N}(y,s)\,\mid\,|y-x|+|t-s|+N^{-1}\leq\delta\right\}$
satisfies $\partial_{t}F_{\infty}-\mathcal{F}(\partial_{x}F_{\infty},\partial_{x}^{2}F_{\infty})\geq 0$ in $\mathbb{R}\times(0,\infty)$ and $F_{\infty}$ equals the left-continuous version of $F_{\text{in}}$ at $t=0$. Therefore, by the comparison principle, $\overline{F}\leq F_{\infty}$ in $\mathbb{R}\times(0,\infty)$, and this implies $\overline{F}\leq F_{\infty}\leq\liminf\nolimits_{*}F_{\epsilon}$ as desired. ∎
B.3. Contractivity
This subsection establishes that operators such as the family $(T_{\epsilon})_{\epsilon>0}$ are contractive on $CDF(\overline{\mathbb{R}})$ with respect to the supremum norm.

### v2

@@ B.1. Proof of Theorem 3
(86) $G^{\star}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{G^{(N)}_{n}(y)\,\mid\,|\delta^{(N)}y-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\}.$
Thus, in passing from $F$’s to $G$’s, nothing has changed except the stricter upper bound $G^{(N)}_{n}\leq 1-\delta(1+LN^{-1})^{n}$ holds for each $n$ and the evolution equation has been weakened to an inequality in (85).
Suppose that $\varphi$ is a smooth function defined in some open subset of space-time that touches $G^{\star}$ from above at some point $(x_{0},t_{0})\in\mathbb{R}\times(0,\infty)$ in its domain. The goal is to show that
(87) $\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))+Q(\varphi(x_{0},t_{0}))\leq 0.$
Since $G^{\star}(x,t)\geq\max\{c_{-}(t)-\delta e^{Lt},0\}$ pointwise in $\mathbb{R}\times[0,+\infty)$ by (84) and the definition of $G^{\star}$, it is convenient to consider two cases separately:
$\text{(I)}\quad\varphi(x_{0},t_{0})=\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}\quad\text{and}\quad\text{(II)}\quad\varphi(x_{0},t_{0})>\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}.$
Step 3, Case (I). First, consider case (I). Since $\varphi\geq G^{\star}$ holds in a neighborhood of $(x_{0},t_{0})$, $\varphi(x_{0},t_{0})=G^{*}(x_{0},t_{0})$, and $G^{\star}(x,t)\geq\max\{c_{-}(t)-\delta e^{Lt},0\}$ globally, in this case, $(x_{0},t_{0})$ is a local minimum of the function $(x,t)\mapsto\varphi(x,t)-\max\{c_{-}(t)-\delta e^{Lt},0\}$. In particular, $x_{0}$ is a local minimum of $x\mapsto\varphi(x,t_{0})$, hence
$\partial_{x}\varphi(x_{0},t_{0})=0,\quad\partial_{x}^{2}\varphi(x_{0},t_{0})\geq 0.$
Therefore, since $\mathcal{F}(0,0)=0$ by (48) and $\mathcal{F}$ is nondecreasing in the second argument,
(88) $\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))=\mathcal{F}(0,\partial_{x}^{2}\varphi(x_{0},t_{0}))\geq 0.$
Restricting attention this time to the slice $x=x_{0}$, the function $t\mapsto\varphi(x_{0},t)-\max\{c_{-}(t)-\delta e^{Lt},0\}$ has a local minimum at $t=t_{0}$, from which it readily follows that either
$\partial_{t}\varphi(x_{0},t_{0})=\frac{d}{dt}\{c_{-}(t)-\delta e^{Lt}\}=-Q(c_{-}(t_{0}))-L\delta e^{Lt_{0}}\leq-Q(\varphi(x_{0},t_{0}))$
if $c_{-}(t_{0})-\delta e^{Lt_{0}}>0$, or
$\partial_{t}\varphi(x_{0},t_{0})=0\leq-Q(0)=-Q(\varphi(x_{0},t_{0}))$
if $c_{-}(t_{0})-\delta e^{Lt_{0}}\leq 0$. (That $Q(0)\leq 0$ is a consequence of the fact that the function $c\mapsto c-Q(c)$ maps $[0,1]$ into itself by assumptions (i) and (iv); see the discussion in Section 5.1.) Combining the inequality $\partial_{t}\varphi(x_{0},t_{0})\leq-Q(\varphi(x_{0},t_{0}))$ with (88), one obtains the desired inequality (87).
Step 3, Case (II). It only remains to consider case (II). That is, from now on, assume that $\varphi(x_{0},t_{0})>\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}$.
According to Lemma 4, there is no loss of generality in assuming that either
$\text{(a)}\quad\partial_{x}\varphi(x_{0},t_{0})\neq 0\quad\text{or}\quad\text{(b)}\quad\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0.$
In case (a), since $G^{\star}$ is nondecreasing, it follows that $\partial_{x}\varphi(x_{0},t_{0})>0$, and Lemma 2 is applicable. In case (b), Lemma 3 applies. To avoid repetition, only case (b) will be considered in what follows.
Therefore, assume henceforth that (b) holds. According to Lemma 3 with $A=1-\frac{\delta}{2}$, there is no loss of generality in fixing an $r>0$ and assuming that
$\varphi(x,t)=\varphi_{1}(x)+\varphi_{2}(t)\quad\text{for each}\,\,(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r],$
where $\varphi_{1}$ is a nondecreasing $C^{2}$ function with $\varphi_{1}(x)=\varphi_{1}(x_{0})$ for $x\leq x_{0}$ and $\varphi_{1}(x)=1-\frac{\delta}{2}$ for $x\geq x_{0}+r$; $\varphi_{2}$ is $C^{2}$ in $t_{0}-r,t_{0}+r]$ with $\varphi_{2}(t_{0})=0$; and
@@
Since $\varphi_{2}$ is continuous in $t_{0}-r,t_{0}+r]$ and $\varphi_{1}(-\infty)=\varphi(x_{0},t_{0})>\max\{c_{-}(t_{0})-\delta e^{Lt_{0}},0\}$, up to shrinking $r$ if necessary, there is no loss of generality in assuming that there are constants $c_{*},c^{*}\in[0,1]$ such that, for each $t\in[t_{0}-r,t_{0}+r]$,
(89) $\max\{c_{-}(t)-\delta e^{Lt},0\}<c_{*}\leq\varphi(-\infty,t)\quad\text{and}\quad 1-\delta<c_{*}\leq\varphi(+\infty,t)\leq 1-\frac{\delta}{4}.$
Finally, for each $N>0$, let $\nu(N)$ be the constant
$\nu(N)=\sup\left\{G^{(N)}_{n}(y)-\varphi(\delta^{(N)}y,N^{-1}n)\,\mid\,y\in\mathbb{R},\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}.$
Notice that (89), (84), and (83) imply that there is an $R>0$ such that, for any $N$,
$\sup\left\{G^{(N)}_{n}(y)-\varphi(\delta^{(N)}y,N^{-1}n)\,\mid\,\delta^{(N)}|y|\geq R,\,\,N^{-1}n\in[t_{0}-r,t_{0}+r]\right\}<0.$
Therefore, by the equation (86) relating $\{G^{(N)}_{n}\}$ to $G^{\star}$, it follows that there is a sequence $\{(N_{k},y_{k},n_{k})\}$ such that $(N^{-1}_{k},\nu(N_{k}))\to(0,0)$ as $k\to\infty$; $\delta^{(N)}|y_{k}|\leq R$ and $|N^{-1}n_{k}-t_{0}|\leq r$ for each $k$; and
(90) $G^{(N_{k})}_{n_{k}}(y_{k})-\varphi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})=\nu(N_{k})\quad\text{for each}\,\,k.$
Further, since the set $\{(x,t)\in\mathbb{R}\times[t_{0}-r,t_{0}+r]\,\mid\,G^{\star}(x,t)=\varphi(x,t)\}$ is contained in $(-\infty,x_{0}]\times\{t_{0}\}$, there is no loss of generality in assuming that there is an $x^{\prime}_{0}\leq x_{0}$ such that $(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k})\to(x^{\prime}_{0},t_{0})$ and $G^{\star}(x^{\prime}_{0},t_{0})=\varphi(x^{\prime}_{0},t_{0})$. Note, in particular, that the inequality $x^{\prime}_{0}\leq x_{0}$ implies, by the constancy of $\varphi_{1}$ in $(-\infty,x_{0}]$,
(91) $\lim_{k\to\infty}(\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\varphi_{1}(\delta^{(N_{k})}y_{k}))=(\varphi_{1}(x_{0}),0,0).$
Finally, for notational ease, let $\Phi_{k}:\mathbb{R}\times[t_{0}-r,t_{0}+r]\to\mathbb{R}$ be the function $\Phi_{k}(y,t)=\varphi(\delta^{(N_{k})}y,t)+\nu(N_{k})$. Since $\nu(N_{k})\to 0$, the bounds (89) imply that $0\leq\Phi_{k}\leq 1$ pointwise for all $k$ large enough. That is, $\Phi_{k}\in CDF(\overline{\mathbb{R}})$ for large enough $k$. By the choice of $\nu(N_{k})$,
$G^{(N_{k})}_{n}\leq\Phi_{k}(\cdot,N_{k}^{-1}n)\quad\text{pointwise for each}\,\,n\in[N(t_{0}-r),N(t_{0}+r)].$
Therefore, by monotonicity (assumption (i)),
$T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}\leq T^{(N_{k})}\Phi_{k}(\cdot,N_{k}^{-1}(n_{k}-1)).$
@@
$\Phi_{k}(y_{k},N_{k}^{-1}n_{k})=G^{(N_{k})}_{n_{k}}(y_{k})\leq T^{(N_{k})}G^{(N_{k})}_{n_{k}-1}(y_{k})\leq T^{(N_{k})}\Phi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1}).$
@@
$N_{k}\Big(\Phi_{k}(y_{k},N_{k}^{-1}n_{k})-\Phi_{k}(y_{k},N_{k}^{-1}n_{k}-N_{k}^{-1})\Big)$
$\quad\quad\quad\leq N_{k}(\mathscr{L}^{(N_{k})}\Phi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))+N_{k}Q^{(N_{k})}(\Phi_{k}(y_{k},N_{k}^{-1}(n_{k}-1)).$
By assumption (iii), the definition of $\Phi_{k}$, and the form of $\varphi$,
$(\mathscr{L}^{(N_{k})}\Phi_{k})(y_{k},N_{k}^{-1}(n_{k}-1))=(\mathscr{L}^{(N_{k})}(\varphi_{1})_{\delta^{(N_{k})}})(y_{k}),$
hence, by assumptions (v) and (vi), (91), and the form of $\varphi$, sending $k\to\infty$ above yields
$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(0,0)+Q(\varphi(x_{0},t_{0})))$ $=\partial_{t}\varphi_{2}(t_{0})-\lim_{k\to\infty}\mathcal{F}(\partial_{x}\varphi_{1}(\delta^{(N_{k})}y_{k}),\partial_{x}^{2}\varphi_{1}(\delta^{(N_{k})}y_{k}))$
$\quad+\lim_{k\to\infty}N_{k}Q^{(N_{k})}(\varphi(\delta^{(N_{k})}y_{k},N_{k}^{-1}n_{k}))$
$\leq 0.$
Since $\partial_{x}\varphi(x_{0},t_{0})=\partial_{x}^{2}\varphi(x_{0},t_{0})=0$ by assumption, this yields (87) as desired.
Step 4: $F_{\star}$ is a Supersolution. The proof that $F_{\star}$ is a supersolution of (61) can be reduced to the subsolution property of $F^{\star}$ after a suitable transformation. As in Remark 2 above, let $G_{-}$ denote the left-continuous version of a nondecreasing function $G$, and define an involution $G\mapsto G_{\text{rev}}$ on $CDF(\overline{\mathbb{R}})$ via the formula
@@
Let $\{\tilde{T}^{(N)}\}$ be the operators on $CDF(\overline{\mathbb{R}})$ obtained from $\{T^{(N)}\}$ by
$\tilde{T}^{(N)}F=(T^{(N)}(F_{\text{rev}}))_{\text{rev}}.$
It is straightforward to check that the family $\{\tilde{T}^{(N)}\}$ satisfies assumptions (i)-(vi), the only difference being $\mathcal{F}$ has to be replaced by the function $\tilde{\mathcal{F}}(v,w)=-\mathcal{F}(v,-w)$ and $Q$ by $\tilde{Q}(q)=-Q(1-q)$.
Define $\{\tilde{F}^{(N)}_{n}\}$ by the rule $\tilde{F}^{(N)}_{n}=(F^{(N)}_{n})_{\text{rev}}$. Notice that the recursion $F^{(N)}_{n}=T^{(N)}F^{(N)}_{n-1}$ implies that $\tilde{F}^{(N)}_{n}=\tilde{T}^{(N)}\tilde{F}^{(N)}_{n-1}$. Therefore, by Step 3, if $\tilde{F}^{\star}$ is defined by
$\tilde{F}^{\star}(x,t)=\lim_{\delta\downarrow 0}\sup\left\{\tilde{F}^{(N)}_{n}(y)\,\mid\,|\delta^{(N)}y-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\},$
then $\tilde{F}^{\star}$ is a subsolution of the initial value problem (61) with $\mathcal{F}$ replaced by $\tilde{\mathcal{F}}$, $Q$ by $\tilde{Q}$, and $F_{\text{in}}$ by $(F_{\text{in}})_{\text{rev}}$. Further, by definition, $\tilde{F}^{\star}(x,t)=1-F_{\star}(-x,t)$. From these last two observations, one readily concludes that $F_{\star}$ is a viscosity supersolution of (61).∎
@@ B.2. Proof of Corollary 4
In the proof that follows, it it convenient to work with the half-relaxed limits $\liminf_{*}F_{N}$ and $\limsup_{*}F_{N}$ of the rescaled CDF $F_{N}$ defined by
$(\liminf\nolimits_{*}F_{N})(x,t)$ $=\lim_{\delta\downarrow 0}\inf\left\{F_{N}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|s-t|+N^{-1}\leq\delta\right\},$
$(\limsup\nolimits_{*}F_{N})(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{\epsilon}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|s-t|+N^{-1}\leq\delta\right\}.$
@@
Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. Suppose that, at time $t=0$, $F_{N}(\cdot,0)\to F_{\text{in}}$ vaguely as $N\to+\infty$. Let $\underline{F}$ and $\overline{F}$ be the maximal subsolution and the minimal supersolution of the problem (61). In view of Proposition 13 and well-known properties of half-relaxed limits (see [15, Remark 6.4] or [10, Lemma 6.2]), to see that $F_{N}$ converges locally uniformly in $\mathbb{R}\times(0,+\infty)$ to the discontinuous viscosity solution as $N\to+\infty$, it suffices to establish that
$\underline{F}\leq\liminf\nolimits_{*}F_{\epsilon}\leq\limsup\nolimits^{*}F_{\epsilon}\leq\overline{F}\quad\text{in}\,\,\mathbb{R}\times[0,\infty).$
@@
For each $M\in\mathbb{N}$, generate a rescaled CDF $F_{N,M}$ like $F_{N}$, albeit instead starting from the initial datum
$F_{N,M}(x,0)=2^{M}\int_{0}^{\infty}F_{N}(x-y,0)\rho(2^{M}y)\,dy,$
where $\rho:(0,\infty)\to(0,\infty)$ is a smooth function supported in $[1,2]$ with $\int_{0}^{\infty}\rho(y)\,dy=1$. From the fact that $\rho$ is supported in $[1,2]$ and $F_{N}$ is nondecreasing in $x$, it follows that $F_{N,M}(x,0)\leq F_{N,M+1}(x,0)\leq F_{N}(x,0)$ for every $x\in\mathbb{R}$. Due to the monotonicity assumption (i), this implies that $F_{N,M}\leq F_{N,M+1}\leq F_{N}$ in $\mathbb{R}\times[0,\infty)$ for each $N$ and $M$. By Theorem 3, $F_{N,M}\to F_{M}$ locally uniformly in $\mathbb{R}\times[0,+\infty)$ as $N\to+\infty$, where $F_{M}$ is the solution of (61) with initial datum $F_{\text{in},M}$ given by $F_{\text{in},M}(x)=2^{M}\int_{0}^{\infty}F_{\text{in}}(x-y)\rho(2^{M}y)\,dy$. Further, by construction, the bound $F_{M}\leq F_{M+1}\leq\liminf\nolimits_{*}F_{N}$ holds pointwise in $\mathbb{R}\times[0,\infty)$ for any fixed $M$.
Standard arguments (see [10, Theorem 6.2]) show that the function $F_{\infty}$ defined by
$F_{\infty}(x,t)=\lim_{\delta\downarrow 0}\inf\left\{F_{M}(x^{\prime},s)\,\mid\,|x^{\prime}-x|+|t-s|+M^{-1}\leq\delta\right\}$
satisfies $\partial_{t}F_{\infty}-\mathcal{F}(\partial_{x}F_{\infty},\partial_{x}^{2}F_{\infty})+Q(F_{\infty})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and $F_{\infty}(\cdot,0)\geq F_{M}(\cdot,0)$ for any $M$. In particular, by the comparison principle, $F_{\infty}$ is at least as large as any viscosity subsolution of the initial-value problem (61). Therefore, $F_{\infty}\geq\underline{F}$ in $\mathbb{R}\times[0,\infty)$, and this implies $\liminf\nolimits_{*}F_{N}\geq F_{\infty}\geq\underline{F}$ as desired. ∎
B.3. Lipschitz Estimate
This subsection establishes that operators such as the family $\{T^{(N)}\}$ are uniformly Lipschitz on $CDF(\overline{\mathbb{R}})$ with the supremum norm.